In 1997, an article on Americans' working habits concluded that the overall proportion of American that drank coffee before work in the morning was .46. As a coffee junkie, you believe that the proportion of all American who drink coffee in the morning has increased since then. Therefore, you take a random sample of 36 Americans and find that 18 of them, or .50, report drinking coffee in the morning before work. Provide a confidence interval for the true proportion of Americans who drink coffee in the morning before work.

Two-sample proportion

Two dependent means

One-sample proportion

One-sample mean

Two independent means

cannot do, assumptions violated
Answer #1

Answer:

One-sample proportion

n=36, p = 0.46, x=18 , = 0.50

in order to do one sample proportion confidence interval the following

assumptions should met.

i) sample should be random

ii)

n * ≥ 10 and n(1- ) ≥ 10

here, sample is random and

n * ≥ 10 = 36 * 0.50 = 18

n(1- ) ≥ 10 = 36 * ( 1 - 0.50 ) = 18

thus both are greater than 10 hence, we can say

assumptions are met.